### v6

@@ 1 Introduction
The other extreme is when just enough cycles are present for one atom to be flexible. This case was handled by Alm, Maddux & Manske [AMM08], who exhibited a representation of the algebra $A_{n}$ obtained from splitting the non-flexible diversity atom of $6_{7}$ into $n$ symmetric atoms. [^2] Here, we use Maddux’s [Mad06] numbering for relation algebras such as $6_{7}$ and $32_{65}$. Note that $6_{7}$ is the relation algebra with two symmetric atoms $a,b$, where $a$ is flexible and the sole forbidden diversity cycle is $bbb$. We refer to Table 1 for $32_{65}$. In particular, this construction yielded a representation of $A_{2}=32_{65}$ over a set of size $416,714,805,91$. Dodd & Hirsch [DH13] subsequently improved the upper bound of the minimum representation size of $32_{65}$ to $63,432,274,896$. This was subsequently improved to 8192 by J.F. Alm & D. Sexton (unpublished), and later $3432$ by [AA19]. Finally, in [ALM+22], the authors exhibited a representation over a set of size $1024$ for $32_{65}$, as well as the first polynomial upper bound on $\min(\operatorname{Spec}(A_{n}))$ (where $\operatorname{Spec}(A_{n})$ is the spectrum taken over all square representations).
In the quest for finite representations, it is desirable to constrain the search space. Furthermore, the minimum number of points required to represent a relation algebra serves as a measure of combinatorial complexity. This simultaneously motivates the study of small representations, the study of highly symmetric representations such as over groups, as well as lower bounds. There are few lower bounds in the literature. Jipsen, Maddux, & Tuza exhibited a lower bound of $n^{2}+n+1$ for the relation algebra $\mathfrak{E}_{n+1}(1,2,3)$ [JMT95]. In [ALM+22], the authors showed that any representation of $A_{n}$ requires at least $12n^{2}+4n+$ 1 points, which is asymptotically double the trivial lower bound of $n^{2}+2n+3$. The key technique involved analyzing the combinatorial substructure induced by the flexible atom. In the special case of $A_{2}=32_{65}$, the authors used a SAT solver to further improve the lower bound. Namely, they showed that $32_{65}$ is not representable on a set of fewer than $26$ points.
@@
There has been considerable progress in constructing Ramsey schemes for a finite number of colors using a technique due to Comer [Com83]. Intuitively, Comer’s method takes a positive integer $m$ and considers finite fields $\mathbb{F}_{q}$, where $q\equiv 1\pmod{2m}$. We next consider the unique multiplicative subgroup $H\leq\mathbb{F}_{q}^{\times}$ of order $(q-1)/m$, and check whether the cosets of $\mathbb{F}_{q}^{\times}/H$ yield a group representation (namely, taking the cosets to be the atoms of the relation algebra– see Definition 2.10). With the sole exception of an alternate construction of the $3$-color algebra using $(\mathbb{Z}/4\mathbb{Z})^{2}$ [Whi75], all known constructions have been due to the guess-and-check finite field method of Comer [Com83]. The $m$-color Ramsey number
@@
Alm & Levet [AL23] generalized the notion of a Ramsey scheme to the directed (antisymmetric) setting. As a consequence, they gave finite representations for the relation algebras $33_{37},35_{37}$, $77_{83}$, $78_{83}$, $80_{83}$, $82_{83}$, $83_{83}$, $1310_{1316}$, $1313_{1316}$, $1315_{1316}$, and $1316_{1316}$. Only $83_{83}$ and $1316_{1316}$ were previously known to be finitely representable, by a slight generalization of [JMT95].
@@
As a first application, we provide the first finite representation for the relation algebra $34_{65}$. This relation algebra has four symmetric atoms $1^{\prime},a,b,$ and $c$, with forbidden cycles $bbc$ and $ccb$. The atom $a$ is flexible. Hence, by [Com84, Theorem 5.3], $34_{65}$ admits a representation over a countable set.
Theorem 1.1.
@@ 1.1 Mandatory cycles for algebras used
$aaa$ $bbb$ $ccc$ $abb$ $baa$ $acc$ $caa$ $bcc$ $cbb$ $abc$
$32_{65}$ $aaa$ $abb$ $baa$ $acc$ $caa$ $abc$
$33_{65}$ $aaa$ $bbb$ $abb$ $baa$ $acc$ $caa$ $abc$
$34_{65}$ $aaa$ $bbb$ $ccc$ $abb$ $baa$ $acc$ $caa$ $abc$
$42_{65}$ $aaa$ $bbb$ $ccc$ $abb$ $caa$ $bcc$ $abc$
@@ 2 Preliminaries
  • $\langle A,\leq,\land,\vee,\neg,0,1\rangle$ is a Boolean algebra, with $\neg$ the unary negation operator, 00 the identity for $\vee$, and $1$ the identity for $\land$.
@@
  • $\breve{}$   is the unary converse operation, and the reduct $\langle A,\circ,\breve{}\;\rangle$ is an involuted monoid. Namely, $\breve{\breve{a}}=a$ for all $a\in A$, and
@@
Definition 2.5.
For atoms $a,b,c$, the triple $(a,b,c)$ – usually denoted $abc$ – is called a cycle. We say that the cycle $abc$ is forbidden if $(a\circ b)\land c=0$ and mandatory if $a\circ b\geq c$. If $a,b,c$ are diversity atoms, then $abc$ is called a diversity cycle.
@@
Definition 2.7.
Let $f$ be a symmetric diversity atom. We say that $f$ is flexible if for all diversity atoms $a,b$, we have that $abf$ is mandatory.
Definition 2.8.
@@
$\text{Spec}(A):=\{\alpha\leq\omega:A\text{ has a square representation over a set of cardinality }\alpha\}.$
@@
:${}^{-1}:2^{G}\to 2^{G}$ maps $X\mapsto X^{-1}:=\{x^{-1}:x\in X\}$, and $e$ is the identity in the group. A group representation of the relation algebra $A$ over the group $G$ is an injective homomorphism $\varphi:A\to\text{Cm}(G)$.
Remark 2.11.
@@
Given a complete graph on $m$ vertices, we seek to color the edges using the diversity atoms as colors. Now suppose $a\in A$ is a diversity atom, and $uv$ is an edge colored $a$ (we abuse notation by using the diversity atom as the label for the edge color). If $abc$ is a mandatory diversity cycle, then there must exist a vertex $w$ such that $vw$ is colored $b$ and $uw$ is colored $c$. If instead $abc$ is a forbidden diversity cycle, then for any vertex $w$, we have that either $vw$ is not colored $b$ or $uw$ is not colored $c$. Now if there exists such a finite $m$, then $A$ is finitely representable. We may view the representation size $m$ as a measure of combinatorial complexity for a relation algebra, and so it is of interest to find the smallest such $m$.
@@ 2.1 Ramsey Schemes
The usual method of constructing the relations $R_{0},\ldots,R_{m-1}$ is a guess-and-check approach due to Comer [Com83], which works as follows. Fix $m\in\mathbb{Z}^{+}$. We search over primes $p\equiv 1\pmod{2m}$, where a desirable $p$ satisfies the following. Let $X_{0}:=H\leq\mathbb{F}_{p}^{\times}$ be the unique subgroup of order $(p-1)/m$. Now let $X_{1},\ldots,X_{m-1}$ be the cosets of $\mathbb{F}_{p}^{\times}/X_{0}$. In particular, as $\mathbb{F}_{p}^{\times}$ is cyclic, we may write $X_{i}=g^{i}X_{0}=\{g^{am+i}:a\in\mathbb{Z}^{+}\},$ where $g$ is a generator of $\mathbb{F}_{p}^{\times}$. Suppose that $X_{0},\ldots,X_{m-1}$ satisfy the following conditions:
@@
    For all distinct $10\leq i,j\leq m-$ 1, $X_{i}+X_{j}=\mathbb{F}_{p}\setminus\{0\}$.
For each $10\leq i\leq m-$ 1, define $R_{i}:=\{(x,y)\in\mathbb{F}_{p}\times\mathbb{F}_{p}:x-y\in X_{i}\}.$ Here, the sets $R_{0},\ldots,R_{m-1}$ together with $\text{Id}=\{(u,u):u\in\mathbb{F}_{p}\}$ are the atoms in our relation algebra. It is easy to check that conditions (a)–(c) on the sets $X_{0},\ldots,X_{m-1}$ imply that conditions (A)–(C) from the definition of a Ramsey scheme are satisfied for the relations $R_{0},\ldots,R_{m-1}$.
@@ 3 Integral Symmetric Relation Algebras with Forbidden Cycle Configurations
Definition 3.1.
Let $p$ be a prime, $U=\mathbb{F}_{p}$, and $m$ a divisor of $(p-1)/2$. Let $H\leq\mathbb{F}_{p}^{\times}$ be the unique subgroup of index $m$, and let $X_{0},\ldots,X_{m-1}$ be the cosets of $\mathbb{F}_{p}^{\times}/H$. A Comer scheme in $m$ colors is a partition of $U\times U$ into $m+1$ binary relations $\text{Id},R_{0},\ldots,R_{m-1}$ as follows. Let $\text{Id}=\{(x,x):x\in\mathbb{F}_{p}\}$ and $R_{i}=\{(x,y)\in\mathbb{F}_{p}\times\mathbb{F}_{p}:x-y\in X_{i}\}$.
Remark 3.2.
@@
We construct Comer schemes using Comer’s method [Com83]. Fix $m\in\mathbb{Z}^{+}$, and let $p\equiv 1\pmod{2m}$ be a prime. Let $X_{0}:=H\leq\mathbb{F}_{p}^{\times}$ be a subgroup of order $(p-1)/m$. Now let $X_{1},\ldots,X_{m-1}$ be the cosets of $\mathbb{F}_{p}^{\times}/X_{0}$. In particular, as $\mathbb{F}_{p}^{\times}$ is cyclic, we may write $X_{i}=g^{i}X_{0}=\{g^{am+i}:a\in\mathbb{Z}^{+}\},$ where $g$ is a generator of $\mathbb{F}_{p}^{\times}$. Suppose that $X_{0},\ldots,X_{m-1}$ satisfy the following conditions:
@@
    For all distinct $10\leq i,j\leq m-$ 1, $X_{i}+X_{j}=\mathbb{F}_{p}\setminus\{0\}$.
For each $10\leq i\leq m-$ 1, define $R_{i}:=\{(x,y)\in\mathbb{F}_{p}\times\mathbb{F}_{p}:x-y\in X_{i}\}.$ Here, the sets $R_{0},\ldots,R_{m-1}$ are the atoms in our relation algebra. It is easy to check that conditions (a) and (b) on the sets $X_{0},\ldots,X_{m-1}$ imply that conditions (A) and (C) from the definition of a Ramsey scheme are satisfied for the relations $R_{0},\ldots,R_{m-1}$. That is, we drop condition (B) from the definition of a Ramsey scheme.
@@
Proof.
Multiply through by $g^{i}$. ∎
@@
Proof.
Suppose $\pi\in\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+1]))$ and $\pi(0)=x$. Since $a_{0}a_{0}a_{1}$ is forbidden, $a_{\pi(0)}a_{\pi(0)}a_{\pi(1)}$ is forbidden as well, so $a_{x}a_{x}a_{\pi(1)}$ is forbidden. But this forces $\pi(1)=x+1\pmod{n}$, since if $xxy$ is forbidden, $y$ must be $x+1$. Similarly, $\pi(2)$ must be $x+2\pmod{n}$, and so on. So $\pi$ must take the form $\pi(s)=x+s\pmod{n}$. All such permutations are clearly in $\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+1]))$, and we have shown they are the only ones. So $\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+1]))\cong\mathbb{Z}/n\mathbb{Z}$. ∎
Lemma 3.8.
@@
Proof.
Let $\rho:\mathrm{At}(\mathcal{A}_{n}([i,i,i+1]))\to\mathrm{At}(\mathcal{A}_{n}([i,i,i+j]))$ be given by $a_{i}\mapsto a_{j\cdot i\pmod{n}}$. Since $\gcd(j,n)=1$, $\rho$ is a bijection. It is easy to check that $\rho$ preserves the forbidden cycles. ∎
@@
Lemma 3.9.
Let $[X_{i},X_{j},X_{\ell}]$ denote the set of cycles of the form $a_{i}a_{j}a_{\ell}$, where the indices are taken modulo $n$. If $C(p,n)$ has forbidden cycles $[X_{i},X_{i},X_{i+j}]$ and $\gcd(j,n)=1$, then $X_{j}$ contains a primitive root $g$, and reindexing using $g$ as a generator will give forbidden cycles $[X_{i},X_{i},X_{i+1}]$.
Proof.
Let $g$ be the primitive root that gives the indexing with forbidden cycles $[X_{i},X_{i},X_{i+j}]$. Now $g^{\ell}$ is also a primitive root modulo $p$ if $\gcd(\ell,p-1)=1$. We want $g^{\ell}\in X_{j}$, so we want to find an integer $a$ with $\gcd(an+j,p-1)=1$. Since $\gcd(j,n)=1$, Dirichlet’s theorem on primes in arithmetic progressions gives some prime $p^{\prime}=an+j$, and clearly $\gcd(p^{\prime},p-1)=1$. Then $g^{p^{\prime}}$ is a primitive root and is in $X_{j}$. ∎
Lemma 3.10.
@@
Proof.
[…]
Example 3.11.
Consider $\mathcal{A}_{6}([i,i,i+2])$. Then $j=x=2$. The permutation (0 2 4)024(0\ 2\ 4)( 0 2 4 ) permutes the forbidden cycles $a_{0}a_{0}a_{2}$, $a_{2}a_{2}a_{4}$, and $a_{4}a_{4}a_{0}$. See Figure 1.
Figure 1: Depiction of the action of the permutation (0 2 4)024(0\ 2\ 4)( 0 2 4 ) on the forbidden cycles of $\mathcal{A}_{6}([i,i,i+2])$.
@@
Proof.
We need to show $(X_{i}+X_{j})\cap X_{0}\neq\varnothing$ for arbitrary $i$ and $j$. Set $q=p$, $A=X_{0}$, $B=X_{i}$, and $C=X_{j}$ in Lemma 3.12. Then $|A|=|B|=|C|=(p-1)/n$. Then we need $|B||C|\geq q^{3}/d^{2}$, which translates to $(p-1)^{4}\geq n^{4}p^{3}$, which is satisfied when $p>n^{4}+5$. Then all diversity cycles are mandatory by Lemma 3.3. ∎

### v7

@@ 1 Introduction
The other extreme is when just enough cycles are present for one atom to be flexible. This case was handled by Alm, Maddux & Manske [AMM08], who exhibited a representation of the algebra $A_{n}$ obtained from splitting the non-flexible diversity atom of $6_{7}$ into $n$ symmetric atoms. [^2] Here, we use Maddux’s [Mad06] numbering for relation algebras such as $6_{7}$ and $32_{65}$. Note that $6_{7}$ is the relation algebra with two symmetric atoms $a,b$, where $a$ is flexible and the sole forbidden diversity cycle is bbbbbb. We refer to Table 1 for $32_{65}$. In particular, this construction yielded a representation of $A_{2}=32_{65}$ over a set of size $416,714,805,91$. Dodd & Hirsch [DH13] subsequently improved the upper bound of the minimum representation size of $32_{65}$ to $63,432,274,896$. This was subsequently improved to 8192 by J.F. Alm & D. Sexton (unpublished), and later $3432$ by [AA19]. Finally, in [ALM+22], the authors exhibited a representation over a set of size $1024$ for $32_{65}$, as well as the first polynomial upper bound on $\min(\operatorname{Spec}(A_{n}))$ (where $\operatorname{Spec}(A_{n})$ is the spectrum taken over all square representations).
In the quest for finite representations, it is desirable to constrain the search space. Furthermore, the minimum number of points required to represent a relation algebra serves as a measure of combinatorial complexity. This simultaneously motivates the study of small representations, the study of highly symmetric representations such as over groups, as well as lower bounds. There are few lower bounds in the literature. Jipsen, Maddux, & Tuza exhibited a lower bound of $n^{2}+n+1$ for the relation algebra $\mathfrak{E}_{n+1}(1,2,3)$ [JMT95]. In [ALM+22], the authors showed that any representation of $A_{n}$ requires at least $2n^{2}+4n+1$ points, which is asymptotically double the trivial lower bound of $n^{2}+2n+3$. The key technique involved analyzing the combinatorial substructure induced by the flexible atom. In the special case of $A_{2}=32_{65}$, the authors used a SAT solver to further improve the lower bound. Namely, they showed that $32_{65}$ is not representable on a set of fewer than $26$ points.
@@
There has been considerable progress in constructing Ramsey schemes for a finite number of colors using a technique due to Comer [Com83]. Intuitively, Comer’s method takes a positive integer $m$ and considers finite fields $\mathbb{F}_{q}$, where $q\equiv 1\pmod{2m}$. We next consider the unique multiplicative subgroup $H\leq\mathbb{F}_{q}^{\times}$ of order (q1)/m(q-1)/m, and check whether the cosets of $\mathbb{F}_{q}^{\times}/H$ yield a group representation (namely, taking the cosets to be the atoms of the relation algebra– see Definition 2.10). With the sole exception of an alternate construction of the $3$-color algebra using $(\mathbb{Z}/4\mathbb{Z})^{2}$ [Whi75], all known constructions have been due to the guess-and-check finite field method of Comer [Com83]. The $m$-color Ramsey number
@@
Alm & Levet [AL23] generalized the notion of a Ramsey scheme to the directed (antisymmetric) setting. As a consequence, they gave finite representations for the relation algebras $33_{37},35_{37}$, $77_{83}$, $78_{83}$, $80_{83}$, $82_{83}$, $83_{83}$, $10_{1316}, 1313$$13_{1316}$, $1315_{1316}$, and $1316_{1316}$. Only $83_{83}$ and $1316_{1316}$ were previously known to be finitely representable, by a slight generalization of [JMT95].
@@
As a first application, we provide the first finite representation for the relation algebra $34_{65}$. This relation algebra has four symmetric atoms $1^{\prime},a,b,$ and $c$, with forbidden cycles bbcbbc and ccbccb. The atom $a$ is flexible. Hence, by [Com84, Theorem 5.3], $34_{65}$ admits a representation over a countable set.
Theorem 1.1
@@ 1.1 Mandatory cycles for algebras used
aaaaaa bbbbbb cccccc abbabb baabaa accacc caacaa bccbcc cbbcbb abcabc
$32_{65}$ aaaaaa abbabb baabaa accacc caacaa abcabc
$33_{65}$ aaaaaa bbbbbb abbabb baabaa accacc caacaa abcabc
$34_{65}$ aaaaaa bbbbbb cccccc abbabb baabaa accacc caacaa abcabc
$42_{65}$ aaaaaa bbbbbb cccccc abbabb caacaa bccbcc abcabc
@@ 2 Preliminaries
  • $\langle A,\leq,\land,\vee,\neg,0,1\rangle$ is a Boolean algebra, with $\neg$ the unary negation operator, 0 the identity for $\vee$, and $1$ the identity for $\land$.
@@
[…]
@@
Definition 2.5
For atoms $a,b,c$, the triple $(a,b,c)$ – usually denoted abcabc – is called a cycle. We say that the cycle abcabc is forbidden if $(a\circ b)\land c=0$ and mandatory if $a\circ b\geq c$. If $a,b,c$ are diversity atoms, then abcabc is called a diversity cycle.
@@
Definition 2.7
Let $f$ be a symmetric diversity atom. We say that $f$ is flexible if for all diversity atoms $a,b$, we have that abfabf is mandatory.
Definition 2.8
@@
[…]
@@
$^{-1}:2^{G}\to 2^{G}$ maps $X\mapsto X^{-1}:=\{x^{-1}:x\in X\}$, and $e$ is the identity in the group. A group representation of the relation algebra $A$ over the group $G$ is an injective homomorphism $\varphi:A\to\text{Cm}(G)$.
Remark 2.11
@@
Given a complete graph on $m$ vertices, we seek to color the edges using the diversity atoms as colors. Now suppose $a\in A$ is a diversity atom, and $uv$ is an edge colored $a$ (we abuse notation by using the diversity atom as the label for the edge color). If abcabc is a mandatory diversity cycle, then there must exist a vertex $w$ such that $vw$ is colored $b$ and $uw$ is colored $c$. If instead abcabc is a forbidden diversity cycle, then for any vertex $w$, we have that either $vw$ is not colored $b$ or $uw$ is not colored $c$. Now if there exists such a finite $m$, then $A$ is finitely representable. We may view the representation size $m$ as a measure of combinatorial complexity for a relation algebra, and so it is of interest to find the smallest such $m$.
@@ 2.1 Ramsey Schemes
The usual method of constructing the relations $R_{0},\ldots,R_{m-1}$ is a guess-and-check approach due to Comer [Com83], which works as follows. Fix $m\in\mathbb{Z}^{+}$. We search over primes $p\equiv 1\pmod{2m}$, where a desirable $p$ satisfies the following. Let $X_{0}:=H\leq\mathbb{F}_{p}^{\times}$ be the unique subgroup of order (p1)/m(p-1)/m. Now let $X_{1},\ldots,X_{m-1}$ be the cosets of $\mathbb{F}_{p}^{\times}/X_{0}$. In particular, as $\mathbb{F}_{p}^{\times}$ is cyclic, we may write $X_{i}=g^{i}X_{0}=\{g^{am+i}:a\in\mathbb{Z}^{+}\},$ where $g$ is a generator of $\mathbb{F}_{p}^{\times}$. Suppose that $X_{0},\ldots,X_{m-1}$ satisfy the following conditions:
@@
    For all distinct $0\leq i,j\leq m-1$, $X_{i}+X_{j}=\mathbb{F}_{p}\setminus\{0\}$.
For each $0\leq i\leq m-1$, define $R_{i}:=\{(x,y)\in\mathbb{F}_{p}\times\mathbb{F}_{p}:x-y\in X_{i}\}.$ Here, the sets $R_{0},\ldots,R_{m-1}$ together with $\text{Id}=\{(u,u):u\in\mathbb{F}_{p}\}$ are the atoms in our relation algebra. It is easy to check that conditions (a)–(c) on the sets $X_{0},\ldots,X_{m-1}$ imply that conditions (A)–(C) from the definition of a Ramsey scheme are satisfied for the relations $R_{0},\ldots,R_{m-1}$.
@@ 3 Integral Symmetric Relation Algebras with Forbidden Cycle Configurations
Definition 3.1
Let $p$ be a prime, $U=\mathbb{F}_{p}$, and $m$ a divisor of (p1)/2(p-1)/2. Let $H\leq\mathbb{F}_{p}^{\times}$ be the unique subgroup of index $m$, and let $X_{0},\ldots,X_{m-1}$ be the cosets of $\mathbb{F}_{p}^{\times}/H$. A Comer scheme in $m$ colors is a partition of $U\times U$ into $m+1$ binary relations $\text{Id},R_{0},\ldots,R_{m-1}$ as follows. Let $\text{Id}=\{(x,x):x\in\mathbb{F}_{p}\}$ and $R_{i}=\{(x,y)\in\mathbb{F}_{p}\times\mathbb{F}_{p}:x-y\in X_{i}\}$.
Remark 3.2
@@
We construct Comer schemes using Comer’s method [Com83]. Fix $m\in\mathbb{Z}^{+}$, and let $p\equiv 1\pmod{2m}$ be a prime. Let $X_{0}:=H\leq\mathbb{F}_{p}^{\times}$ be a subgroup of order (p1)/m(p-1)/m. Now let $X_{1},\ldots,X_{m-1}$ be the cosets of $\mathbb{F}_{p}^{\times}/X_{0}$. In particular, as $\mathbb{F}_{p}^{\times}$ is cyclic, we may write $X_{i}=g^{i}X_{0}=\{g^{am+i}:a\in\mathbb{Z}^{+}\},$ where $g$ is a generator of $\mathbb{F}_{p}^{\times}$. Suppose that $X_{0},\ldots,X_{m-1}$ satisfy the following conditions:
@@
    For all distinct $0\leq i,j\leq m-1$, $X_{i}+X_{j}=\mathbb{F}_{p}\setminus\{0\}$.
For each $0\leq i\leq m-1$, define $R_{i}:=\{(x,y)\in\mathbb{F}_{p}\times\mathbb{F}_{p}:x-y\in X_{i}\}.$ Here, the sets $R_{0},\ldots,R_{m-1}$ are the atoms in our relation algebra. It is easy to check that conditions (a) and (b) on the sets $X_{0},\ldots,X_{m-1}$ imply that conditions (A) and (C) from the definition of a Ramsey scheme are satisfied for the relations $R_{0},\ldots,R_{m-1}$. That is, we drop condition (B) from the definition of a Ramsey scheme.
@@
{shortproof}
Multiply through by $g^{i}$.
@@
Proof 3.8
Suppose $\pi\in\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+1]))$ and $\pi(0)=x$. Since $a_{0}a_{0}a_{1}$ is forbidden, $a_{\pi(0)}a_{\pi(0)}a_{\pi(1)}$ is forbidden as well, so $a_{x}a_{x}a_{\pi(1)}$ is forbidden. But this forces $\pi(1)=x+1\pmod{n}$, since if xxyxxy is forbidden, $y$ must be $x+1$. Similarly, $\pi(2)$ must be $x+2\pmod{n}$, and so on. So $\pi$ must take the form $\pi(s)=x+s\pmod{n}$. All such permutations are clearly in $\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+1]))$, and we have shown they are the only ones. So $\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+1]))\cong\mathbb{Z}/n\mathbb{Z}$.
Lemma 3.9
@@
Proof 3.10
Let $\rho:\mathrm{At}(\mathcal{A}_{n}([i,i,i+1]))\to\mathrm{At}(\mathcal{A}_{n}([i,i,i+j]))$ be given by $a_{i}\mapsto a_{j\cdot i\pmod{n}}$. Since $\gcd(j,n)=1$, $\rho$ is a bijection. It is easy to check that $\rho$ preserves the forbidden cycles.
@@
Lemma 3.11
Let $X_{i},X_{j},X_{\ell}]$ denote the set of cycles of the form $a_{i}a_{j}a_{\ell}$, where the indices are taken modulo $n$. If $C(p,n)$ has forbidden cycles $X_{i},X_{i},X_{i+j}]$ and $\gcd(j,n)=1$, then $X_{j}$ contains a primitive root $g$, and reindexing using $g$ as a generator will give forbidden cycles $X_{i},X_{i},X_{i+1}]$.
Proof 3.12
Let $g$ be the primitive root that gives the indexing with forbidden cycles $X_{i},X_{i},X_{i+j}]$. Now $g^{\ell}$ is also a primitive root modulo $p$ if $\gcd(\ell,p-1)=1$. We want $g^{\ell}\in X_{j}$, so we want to find an integer $a$ with $\gcd(an+j,p-1)=1$. Since $\gcd(j,n)=1$, Dirichlet’s theorem on primes in arithmetic progressions gives some prime $p^{\prime}=an+j$, and clearly $\gcd(p^{\prime},p-1)=1$. Then $g^{p^{\prime}}$ is a primitive root and is in $X_{j}$.
Lemma 3.13
@@
Proof 3.14
Let $x=\gcd(j,n)>1$. Consider the permutation $\pi=(0\ x\ 2x\ 3x\ldots)$, written in cycle notation. We claim that $\pi\in\operatorname{Aut}(\mathcal{A}_{n}([i,i,i+j])))$. Consider the forbidden cycle $a_{0}a_{0}a_{j}$. Write $j=bx$ for some positive integer $b$. Under $\pi$, this cycle $a_{0}a_{0}a_{bx}$ gets mapped to $a_{x}a_{x}a_{(b+1)x}$, and since $(b+1)x=x+j$, the cycle $a_{x}a_{x}a_{(b+1)x}$ is forbidden. In fact, $\pi$ just permutes the forbidden cycles $a_{\ell x}a_{\ell x}a_{\ell x+j}$ and leaves the other forbidden cycles fixed.
Example 3.15
Consider $\mathcal{A}_{6}([i,i,i+2])$. Then $j=x=2$. The permutation (0 2 4)(0\ 2\ 4) permutes the forbidden cycles $a_{0}a_{0}a_{2}$, $a_{2}a_{2}a_{4}$, and $a_{4}a_{4}a_{0}$. See Figure 1.
Figure 1: Depiction of the action of the permutation (0 2 4)(0\ 2\ 4) on the forbidden cycles of $\mathcal{A}_{6}([i,i,i+2])$.
@@
Proof 3.18
We need to show $(X_{i}+X_{j})\cap X_{0}\neq\varnothing$ for arbitrary $i$ and $j$. Set $q=p$, $A=X_{0}$, $B=X_{i}$, and $C=X_{j}$ in Lemma 3.16. Then |A|=|B|=|C|=(p1)/n|A|=|B|=|C|=(p-1)/n. Then we need $|B||C|\geq q^{3}/d^{2}$, which translates to $(p-1)^{4}\geq n^{4}p^{3}$, which is satisfied when $p>n^{4}+5$. Then all diversity cycles are mandatory by Lemma 3.3.